OpenStudy (anonymous):

Power rule and natural log 2 approximation determine approximation for ln 8.

OpenStudy (anonymous):

ln8 = ln(2^3) = 3ln2..

OpenStudy (anonymous):

Because: \[\ln(a^b) = b.\ln(a)\]